Many chemical reactions release energy. Suppose that at the beginning of a reaction, an electron and proton are separated by 0.110 nm , and their final separation is 0.104 nm.

How much electric potential energy was lost in this reaction (in units of eV)?
Answer #1

Comment below if you have any query.

Electric potential energy is given by:

PE = -k*q1*q2/r

here, k = electrostatic constant = 9*10^9

q1 = -e = charge on electron = -1.6*10^-19 C

q2 = +e = charge on proton = +1.6*10^-19 C

Now

r1 = initial separation = 0.110 nm = 0.110*10^-9 m

r2 = final separation = 0.104 nm = 0.104*10^-9 m

So Electric potential energy lost will be

dPE = PEf - PEi

dPE = (-k*q1*q2/r2) - (-k*q1*q2/r1)

dPE = k*q1*q2*(1/r1 - 1/r2)

Using known values:

dPE = 9*10^9*(-1.6*10^-19)*(1.6*10^-19)*(1/(0.110*10^-9) - 1/(0.104*10^-9))

dPE = 1.208*10^-19 J = 1.208*10^-19/(1.6*10^-19)

dPE = 0.755 eV

Please Upvote.